OpenStudy (anonymous):

Help simplifying x= t/(t+1)

OpenStudy (anonymous):

\[x=\frac{ t }{ t+1 }\] they simplified it to \[x=\frac{ 1 }{ 1+\frac{ 1 }{ t } }\]

OpenStudy (anonymous):

just having a hard time processing it.

OpenStudy (shamil98):

It's already simplified..

OpenStudy (anonymous):

ok... i just need to see why i can write it in the other form

OpenStudy (shamil98):

You can simplify it to that because you can take t out of the terms..

OpenStudy (anonymous):

i see. lol

OpenStudy (anonymous):

it didn't get simplified. Both top and bottom are multiplied by (1/t)

OpenStudy (shamil98):

\[x = \frac{ t }{ t+1 } = \frac{ t(1) }{ t(1+\frac{ 1 }{ t }) } = \frac{ \cancel{t}1 }{ \cancel{t}(1+\frac{ 1 }{ t }) } = \frac{ 1 }{ 1+\frac{ 1 }{ t } }\]

OpenStudy (anonymous):

awesome, thanks, I have been looking at it for 10 minutes ... maybe its time to step away for awhile.

OpenStudy (anonymous):

now to go finish the rest of the problem.